# references.bib

### v1

@@ 4.2.2 Connecting elementary cells
$\boldsymbol{r}^{\mathrm{A}}_{0j}$ $=\boldsymbol{r}^{\mathrm{B}}_{0j-1}\,,$ (22a)
$\boldsymbol{r}^{\mathrm{A^{\prime}}}_{0j}$ $=\boldsymbol{r}^{\mathrm{B^{\prime}}}_{0j-1}\,.$ (22b)
Then we have to obtain the point $P^{\prime}$ of our new elementary cell, which has to satisfy the Totimorphic condition. In the following, all points are in the same elementary cell $(0j)$, and thus we drop the indices for convenience. The condition for the existence of $\boldsymbol{r}^{\mathrm{P^{\prime}}}$ is that lever and beam with fixed length have to connect to it from point $A^{\prime}$ (given by the first elementary cell) and point $C$:
$\|\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{P^{\prime}}}\|^{2}$ $=\left(\frac{l_{\text{b}}}{2}\right)^{2}\,,$ (23a)
$\|\boldsymbol{r}^{\mathrm{C}}-\boldsymbol{r}^{\mathrm{P^{\prime}}}\|^{2}$ $=\left(\frac{l_{\text{b}}}{2}\right)^{2}\,.$ (23b)
@@
A detailed derivation can be found in the supplemental information.
4.2.3 Conditions for intact elementary cells
@@
The first condition means that springs cannot be stretched further than the beam length. The second (third) conditions check whether the left (right) bottom and top springs cross.
To avoid overstretched springs, we correct $\theta$ by constraining its range. The limits of this interval are obtained by the criterion for the existence of real-values solutions of Eq. 25 (the term below the root has to be $\geq 0$, see supplemental information):
@@
4.2.4 Building a structure
@@
To initialise the lattice as a flat surface, beam angles $\varphi_{ij}$ and $\varphi^{\prime}_{ij}$ have to be set to 00 and lever angles $\theta_{ij}$ and $\theta^{\prime}_{ij}$ to $\frac{\pi}{2}$ $\forall i,j$.
@@ 4.3 Totimorphic lattices in three dimensions
4.3.1 Parametrisation of elementary cells
@@
$\boldsymbol{r}^{\mathrm{A}}$ $=\boldsymbol{o}\,,$ (30a)
$\boldsymbol{r}^{\mathrm{B}}$ $=\boldsymbol{r}^{\mathrm{A}}+l_{\text{b}}\cdot\boldsymbol{\epsilon}^{\varphi_{y},\varphi_{z}}\,,$ (30b)
$\boldsymbol{r}^{\mathrm{P}}$ $=\boldsymbol{r}^{\mathrm{A}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\epsilon}^{\varphi_{y},\varphi_{z}}\,,$ (30c)
$\boldsymbol{r}^{\mathrm{C}}$ $=\boldsymbol{r}^{\mathrm{P}}+\frac{l_{\text{b}}}{2}\cdot R_{y}(\varphi_{y})R_{z}(\varphi_{z})R_{x}(\theta_{x})R_{z}(\theta_{z})\mathbb{1}_{x}\,.$ (30d)
Here, the beam is described by its rotation in the x-y plane (given by $\varphi_{z}{}$, which corresponds to $\varphi$ in the two-dimensional case) and a rotation in x-z plane (a tilting of the unit cell given by $\varphi_{y}$). The orientation of the lever is obtained by a sequence of rotations: first, we rotate the x-direction unit vector $\mathbb{1}_{x}$ by $\theta_{z}$ (in the two-dimensional case, we called this $\theta$), then tilt it in the y-z plane by $\theta_{x}$, and then apply the same transformation (two rotations) used to gain the beam vector from the unit vector. We choose this parametrisation so that the angles of the lever are again defined with respect to the orientation of the beam, as in the two-dimensional case.
@@
$\boldsymbol{r}^{\mathrm{P^{\prime}}}$ $=\boldsymbol{r}^{\mathrm{C}}+\frac{l_{\text{b}}}{2}\cdot R_{y}(\varphi^{\prime}_{y})R_{z}(\varphi^{\prime}_{z})R_{x}(-\theta^{\prime}_{x})R_{z}(-\theta^{\prime}_{z})\mathbb{1}_{x}\,,$ (31a)
$\boldsymbol{r}^{\mathrm{A^{\prime}}}$ $=\boldsymbol{r}^{\mathrm{P^{\prime}}}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\epsilon}^{\varphi^{\prime}_{y},\varphi^{\prime}_{z}}\,,$ (31b)
$\boldsymbol{r}^{\mathrm{B^{\prime}}}$ $=\boldsymbol{r}^{\mathrm{P^{\prime}}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\epsilon}^{\varphi^{\prime}_{y},\varphi^{\prime}_{z}}\,.$ (31c)
4.3.2 Connecting elementary cells
@@
$\boldsymbol{r}^{\mathrm{A}}_{0j}$ $=\boldsymbol{r}^{\mathrm{B}}_{0j-1}\,,$ (32a)
$\boldsymbol{r}^{\mathrm{A^{\prime}}}_{0j}$ $=\boldsymbol{r}^{\mathrm{B^{\prime}}}_{0j-1}\,.$ (32b)
@@
$\boldsymbol{r}_{L}$ $=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}+\frac{l_{\text{b}}}{2}\boldsymbol{\xi}^{\phi,\Theta}\,,$ (33a)
$=\boldsymbol{s}^{-}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\xi}^{\phi,\Theta}\,.$ (33b)
@@
$\boldsymbol{r}^{\mathrm{P^{\prime}}}=\boldsymbol{r}^{\mathrm{A^{\prime}}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\xi}^{\phi,\Theta}\,,$ (35a)
$\boldsymbol{r}^{\mathrm{B^{\prime}}}=\boldsymbol{r}^{\mathrm{A^{\prime}}}+l_{\text{b}}\cdot\boldsymbol{\xi}^{\phi,\Theta}\,.$ (35b)
A detailed derivation of the angles and their constrained intervals can be found in the supplemental information.
4.3.3 Building a structure

### v2

@@ 4.2.2 Connecting elementary cells in two dimensions
$\boldsymbol{r}^{\mathrm{A}}_{0j}$ $=\boldsymbol{r}^{\mathrm{B}}_{0j-1}\,$, (33)
$\boldsymbol{r}^{\mathrm{A^{\prime}}}_{0j}$ $=\boldsymbol{r}^{\mathrm{B^{\prime}}}_{0j-1}\,.$ (34)
Then we have to obtain the point $P^{\prime}$ of our new elementary cell, which has to satisfy the beam and lever constraints. In the following, all points are in the same elementary cell $(0j)$, and thus we drop the indices for convenience. The condition for the existence of $\boldsymbol{r}^{\mathrm{P^{\prime}}}$ is that lever and beam with fixed length have to connect to it from point $A^{\prime}$ (given by the first elementary cell) and point $C$:
$\|\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{P^{\prime}}}\|^{2}$ $=\left(\frac{l_{\text{b}}}{2}\right)^{2}\,$, (35)
$\|\boldsymbol{r}^{\mathrm{C}}-\boldsymbol{r}^{\mathrm{P^{\prime}}}\|^{2}$ $=\left(\frac{l_{\text{b}}}{2}\right)^{2}\,.$ (36)
@@
4.2.3 Proof: derivation of expression for $\mathrm{P^{\prime}}$
$\|\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{P^{\prime}}}\|^{2}$ $=\left(\frac{l_{\text{b}}}{2}\right)^{2}\,$, (39)
$\|\boldsymbol{r}^{\mathrm{C}}-\boldsymbol{r}^{\mathrm{P^{\prime}}}\|^{2}$ $=\left(\frac{l_{\text{b}}}{2}\right)^{2}\,.$ (40)
Expanding both expressions, subtracting Eq. 40 from Eq. 39 and dividing the resulting equation by $2$, we obtain:
$\frac{1}{2}\left(\|\boldsymbol{r}^{\mathrm{A^{\prime}}}\|^{2}-\|\boldsymbol{r}^{\mathrm{C}}\|^{2}\right)-r^{\mathrm{P^{\prime}}}_{y}\left(r^{\mathrm{A^{\prime}}}_{y}-r^{\mathrm{C}}_{y}\right)-r^{\mathrm{P^{\prime}}}_{x}\left(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x}\right)=0\,.$ (41)
Using the notation $\boldsymbol{s}^{-}=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}$, $\boldsymbol{s}^{+}=\boldsymbol{r}^{\mathrm{A^{\prime}}}+\boldsymbol{r}^{\mathrm{C}}$ and $\delta=\frac{1}{2}\left(\|\boldsymbol{r}^{\mathrm{A^{\prime}}}\|^{2}-\|\boldsymbol{r}^{\mathrm{C}}\|^{2}\right)$ as in the main text, this yields Eq. 37. To obtain Eq. 38, we first insert Eq. 37 into Eq. 40 and rewrite it to match the form of a quadratic equation:
$(r^{\mathrm{P^{\prime}}}_{y})^{2}+p\cdot r^{\mathrm{P^{\prime}}}_{y}+q=0\,$, (42)
$\mathrm{with}\ \ \$$p=\frac{2r^{\mathrm{C}}_{x}s^{-}_{y}s^{-}_{x}-2r^{\mathrm{C}}_{y}(s^{-}_{x})^{2}-2\delta s^{-}_{y}}{(s^{-}_{x})^{2}+(s^{-}_{y})^{2}}\,$, (43)
$q=\frac{\delta^{2}-2\delta r^{\mathrm{C}}_{x}s^{-}_{x}+(r^{\mathrm{C}}_{x}s^{-}_{x})^{2}+(r^{\mathrm{C}}_{y}s^{-}_{x})^{2}-(s^{-}_{x})^{2}(\frac{l_{\text{b}}}{2})^{2}}{(s^{-}_{x})^{2}+(s^{-}_{y})^{2}}\,.$ (44)
Writing out all terms in the numerator of $p$ (denoted by $p_{\mathrm{n}}$), it can be simplified to
$p_{\mathrm{n}}=\$$2r^{\mathrm{C}}_{x}s^{-}_{y}s^{-}_{x}-2r^{\mathrm{C}}_{y}(s^{-}_{x})^{2}-2\delta s^{-}_{y}$
$=\$$2r^{\mathrm{C}}_{x}r^{\mathrm{A^{\prime}}}_{x}r^{\mathrm{A^{\prime}}}_{y}-(r^{\mathrm{C}}_{x})^{2}r^{\mathrm{A^{\prime}}}_{y}-r^{\mathrm{C}}_{y}(r^{\mathrm{A^{\prime}}}_{x})^{2}+2r^{\mathrm{C}}_{x}r^{\mathrm{C}}_{y}r^{\mathrm{A^{\prime}}}_{x}-(r^{\mathrm{A^{\prime}}}_{x})^{2}r^{\mathrm{A^{\prime}}}_{y}$
$-(r^{\mathrm{A^{\prime}}}_{y})^{3}+(r^{\mathrm{C}}_{y})^{2}r^{\mathrm{A^{\prime}}}_{y}+(r^{\mathrm{A^{\prime}}}_{y})^{2}r^{\mathrm{C}}_{y}-(r^{\mathrm{C}}_{x})^{2}r^{\mathrm{C}}_{y}-(r^{\mathrm{C}}_{y})^{3}\,.$ (45)
Using the two identities:
$(r^{\mathrm{A^{\prime}}}_{y})^{3}-(r^{\mathrm{A^{\prime}}}_{y})^{2}r^{\mathrm{C}}_{y}+r^{\mathrm{A^{\prime}}}_{y}(r^{\mathrm{C}}_{x})^{2}-r^{\mathrm{A^{\prime}}}_{y}(r^{\mathrm{C}}_{y})^{2}+(r^{\mathrm{C}}_{x})^{2}r^{\mathrm{C}}_{y}+(r^{\mathrm{C}}_{y})^{3}=(r^{\mathrm{A^{\prime}}}_{y}+r^{\mathrm{C}}_{y})\left[(r^{\mathrm{A^{\prime}}}_{y}-r^{\mathrm{C}}_{y})^{2}+(r^{\mathrm{C}}_{x})^{2}\right]\,$, (46)
$r^{\mathrm{C}}_{y}(r^{\mathrm{A^{\prime}}}_{x})^{2}-2r^{\mathrm{A^{\prime}}}_{x}r^{\mathrm{A^{\prime}}}_{y}r^{\mathrm{C}}_{x}-2r^{\mathrm{C}}_{x}r^{\mathrm{C}}_{y}r^{\mathrm{A^{\prime}}}_{x}+(r^{\mathrm{A^{\prime}}}_{x})^{2}r^{\mathrm{A^{\prime}}}_{y}=(r^{\mathrm{A^{\prime}}}_{y}+r^{\mathrm{C}}_{y})\left[(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x})^{2}-(r^{\mathrm{C}}_{x})^{2}\right]\,$, (47)
we arrive at
$p_{\mathrm{n}}=-s^{+}_{y}\left[(s^{-}_{x})^{2}+(s^{-}_{y})^{2}\right]\,$, (48)
and thus
$p=-s^{+}_{y}\,$. (49)
For $q$, we first note (using binomial formulas) that
$\delta-r^{\mathrm{C}}_{x}s^{-}_{x}=\frac{1}{2}\left[(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x})^{2}+(r^{\mathrm{A^{\prime}}}_{y}-r^{\mathrm{C}}_{y})(r^{\mathrm{A^{\prime}}}_{y}+r^{\mathrm{C}}_{y})\right]\,.$ (50)
Using $\delta^{2}-2\delta r^{\mathrm{C}}_{x}s^{-}_{x}+(r^{\mathrm{C}}_{x}s^{-}_{x})^{2}=(\delta-r^{\mathrm{C}}_{x}s^{-}_{x})^{2}$ as well as $(\frac{p}{2})^{2}=\frac{1}{4}(r^{\mathrm{A^{\prime}}}_{y}+r^{\mathrm{C}}_{y})^{2}\left[(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x})^{2}+(r^{\mathrm{A^{\prime}}}_{y}-r^{\mathrm{C}}_{y})^{2}\right]$, we can then calculate the remaining term $(\frac{p}{2})^{2}-q$ required for solving Eq. 42. First. we only look at all terms in $(\frac{p}{2})^{2}-q$ not containing $l_{\text{b}}$. The numerator of this term is given by:
$-\frac{1}{4}\left[(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x})^{2}+(r^{\mathrm{A^{\prime}}}_{y}-r^{\mathrm{C}}_{y})(r^{\mathrm{A^{\prime}}}_{y}+r^{\mathrm{C}}_{y})\right]^{2}-(r^{\mathrm{C}}_{y})^{2}(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x})^{2}+\frac{1}{4}(r^{\mathrm{A^{\prime}}}_{y}+r^{\mathrm{C}}_{y})^{2}\left[(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x})^{2}+(r^{\mathrm{A^{\prime}}}_{y}-r^{\mathrm{C}}_{y})^{2}\right]\,$,
$=\$$-\frac{1}{4}(s^{-}_{x})^{2}\left[(r^{\mathrm{A^{\prime}}}_{x}-r^{\mathrm{C}}_{x})^{2}+2((r^{\mathrm{A^{\prime}}}_{y})^{2}-(r^{\mathrm{C}}_{y})^{2})-(r^{\mathrm{A^{\prime}}}_{y}+r^{\mathrm{C}}_{y})^{2}+4(r^{\mathrm{C}}_{y})^{2}\right]\,$, (51)
$=\$$-\frac{1}{4}(s^{-}_{x})^{2}\left[(s^{-}_{x})^{2}+(s^{-}_{y})^{2}\right]\,.$ (52)
The terms containing $l_{\text{b}}$ simplify to $\frac{l_{\text{b}}^{2}(s^{-}_{x})^{2}}{4\|\boldsymbol{s}^{-}\|^{2}}$. Using the previous results, we arrive at:
$\left(\frac{p}{2}\right)^{2}-q=\frac{1}{4}(s^{-}_{x})^{2}\cdot\left(\left(\frac{l_{\text{b}}}{\|\boldsymbol{s}^{-}\|}\right)^{2}-1\right)\,.$ (53)
[…]
$r^{\mathrm{P^{\prime}}}_{y}=\frac{1}{2}\left(s^{+}_{y}\pm s^{-}_{x}\sqrt{\left(\frac{l_{\text{b}}}{\|\boldsymbol{s}^{-}\|}\right)^{2}-1}\right)\,.$ (54)
4.2.4 Conditions for intact elementary cells in two dimensions
@@
The first condition means that springs cannot be stretched further than the beam length (for a solution of $\boldsymbol{r}^{\mathrm{P^{\prime}}}$ to exist, Eq. 38 has to be a real number, meaning that the argument of the square root has to be larger or equal to zero. This immediately yields Eq. 55). The second (third) conditions check whether the left (right) bottom and top springs cross.
To avoid overstretched springs, we correct $\theta$ by constraining its range. The limits of this interval are obtained by the criterion for the existence of real-valued solutions of Eq. 38 (the term below the root has to be $\geq 0$):
@@
4.2.5 Proof: deriving allowed range of $\theta$
First, we write out $\boldsymbol{r}^{\mathrm{C}}$ in $\boldsymbol{s}^{-}$:
$\boldsymbol{s}^{-}$ $=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}\,$, (59)
$=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{A}}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\,$, (60)
$=\boldsymbol{k}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\,$, (61)
with $\boldsymbol{k}=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{A}}-\frac{l_{\text{b}}}{2}\boldsymbol{e}^{\varphi}$. This way, we split up $\boldsymbol{s}^{-}$ into two terms – one depending on $\theta$, and one that is independent of $\theta$, summarising the geometric constraints introduced by the previous unit cell. By representing $\boldsymbol{k}$ in polar coordinates, we arrive at
$\boldsymbol{s}^{-}=\|\boldsymbol{k}\|\boldsymbol{e}^{\Phi}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\,$, (62)
with $\Phi=E\left(\boldsymbol{k}\right)$, which simplifies the following derivation of the range of $\theta$. To guarantee that a solution for point $\mathrm{P^{\prime}}$ exists, Eq. 55 has to be fulfilled. By squaring Eq. 55 and using Eq. 62, we get
$\left\|\|\boldsymbol{k}\|\boldsymbol{e}^{\Phi}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\right\|^{2}\leq l_{\text{b}}^{2}\hskip 11.38109pt$ $\Leftrightarrow\hskip 11.38109pt\|\boldsymbol{k}\|^{2}-l_{\text{b}}\|\boldsymbol{k}\|\boldsymbol{e}^{\Phi}\cdot\boldsymbol{e}^{\varphi+\theta}+\frac{1}{4}l_{\text{b}}^{2}\leq l_{\text{b}}^{2}\,$, (63)
$\Leftrightarrow\hskip 11.38109pt\boldsymbol{e}^{\Phi}\cdot\boldsymbol{e}^{\varphi+\theta}\geq\frac{\|\boldsymbol{k}\|}{l_{\text{b}}}-\frac{3}{4}\frac{l_{\text{b}}}{\|\boldsymbol{k}\|}\,$, (64)
$\Leftrightarrow\hskip 11.38109pt\cos\left(\Phi-\varphi-\theta\right)\geq\frac{\|\boldsymbol{k}\|}{l_{\text{b}}}-\frac{3}{4}\frac{l_{\text{b}}}{\|\boldsymbol{k}\|}\,$, (65)
$\Leftrightarrow\hskip 11.38109pt\cos\left(\Delta-\theta\right)\geq C\,$, (66)
$\pm\left(\Delta-\theta\right)\leq\arccos{C}\,$, (67)
from which we get the upper (choosing -) and lower (choosing ++) bound of $\theta$ shown in Eq. 58.
4.2.6 Building a two-dimensional Totimorphic structure
@@
To initialise the lattice as a flat surface, beam angles $\varphi_{ij}$ and $\varphi^{\prime}_{ij}$ have to be set to 0 and lever angles $\theta_{ij}$ and $\theta^{\prime}_{ij}$ to $\frac{\pi}{2}$ $\forall i,j$.
@@ 4.3 Totimorphic lattices in three dimensions
4.3.1 Parametrisation of elementary cells.
@@
$\boldsymbol{r}^{\mathrm{A}}$ $=\boldsymbol{o}\,$, (68)
$\boldsymbol{r}^{\mathrm{B}}$ $=\boldsymbol{r}^{\mathrm{A}}+l_{\text{b}}\cdot\boldsymbol{\epsilon}^{\varphi_{y},\varphi_{z}}\,$, (69)
$\boldsymbol{r}^{\mathrm{P}}$ $=\boldsymbol{r}^{\mathrm{A}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\epsilon}^{\varphi_{y},\varphi_{z}}\,$, (70)
$\boldsymbol{r}^{\mathrm{C}}$ $=\boldsymbol{r}^{\mathrm{P}}+\frac{l_{\text{b}}}{2}\cdot R_{y}(\varphi_{y})R_{z}(\varphi_{z})R_{x}(\theta_{x})R_{z}(\theta_{z})\boldsymbol{1}_{x}\,.$ (71)
Here, the beam is described by its rotation in the x-y plane (given by $\varphi_{z}{}$, which corresponds to $\varphi$ in the two-dimensional case) and a rotation in x-z plane (a tilting of the unit cell given by $\varphi_{y}$). The orientation of the lever is obtained by a sequence of rotations: first, we rotate the x-direction unit vector $\boldsymbol{1}_{x}$ by $\theta_{z}$ (in the two-dimensional case, we called this $\theta$), then tilt it in the y-z plane by $\theta_{x}$, and then apply the same transformation (two rotations) used to gain the beam vector from the unit vector. We choose this parametrisation so that the angles of the lever are again defined with respect to the orientation of the beam, as in the two-dimensional case.
@@
$\boldsymbol{r}^{\mathrm{P^{\prime}}}$ $=\boldsymbol{r}^{\mathrm{C}}+\frac{l_{\text{b}}}{2}\cdot R_{y}(\varphi^{\prime}_{y})R_{z}(\varphi^{\prime}_{z})R_{x}(-\theta^{\prime}_{x})R_{z}(-\theta^{\prime}_{z})\boldsymbol{1}_{x}\,$, (72)
$\boldsymbol{r}^{\mathrm{A^{\prime}}}$ $=\boldsymbol{r}^{\mathrm{P^{\prime}}}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\epsilon}^{\varphi^{\prime}_{y},\varphi^{\prime}_{z}}\,$, (73)
$\boldsymbol{r}^{\mathrm{B^{\prime}}}$ $=\boldsymbol{r}^{\mathrm{P^{\prime}}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\epsilon}^{\varphi^{\prime}_{y},\varphi^{\prime}_{z}}\,.$ (74)
4.3.2 Connecting elementary cells in three dimensions
@@
$\boldsymbol{r}^{\mathrm{A}}_{0j}$ $=\boldsymbol{r}^{\mathrm{B}}_{0j-1}\,$, (75)
$\boldsymbol{r}^{\mathrm{A^{\prime}}}_{0j}$ $=\boldsymbol{r}^{\mathrm{B^{\prime}}}_{0j-1}\,.$ (76)
@@
$\boldsymbol{r}_{L}$ $=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}+\frac{l_{\text{b}}}{2}\boldsymbol{\xi}^{\phi,\Theta}\,$, (77)
$=\boldsymbol{s}^{-}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\xi}^{\phi,\Theta}\,.$ (78)
@@
$\boldsymbol{r}^{\mathrm{P^{\prime}}}=\boldsymbol{r}^{\mathrm{A^{\prime}}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\xi}^{\phi,\Theta}\,$, (80)
$\boldsymbol{r}^{\mathrm{B^{\prime}}}=\boldsymbol{r}^{\mathrm{A^{\prime}}}+l_{\text{b}}\cdot\boldsymbol{\xi}^{\phi,\Theta}\,.$ (81)
4.3.3 Proof: deriving expressions for $\phi$ and $\Theta$
Expanding Eq. 79, we get
$\boldsymbol{s}^{-}\cdot\boldsymbol{\xi}^{\phi,\Theta}=-\frac{\|\boldsymbol{s}^{-}\|^{2}}{l_{\text{b}}}\,.$ (82)
$\boldsymbol{\xi}^{\alpha,\beta}\cdot\boldsymbol{\xi}^{\phi,\Theta}=-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\,$, (83)
or, using Eq. 16,
$\boldsymbol{\xi}^{\alpha,\beta}\cdot\boldsymbol{\xi}^{\phi,\Theta}=\cos\left(\alpha-\phi\right)\sin{\beta}\sin{\Theta}+\cos{\beta}\cos{\Theta}=-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\,$, (84)
from which we get the solution for $\phi$ depending on $\Theta$:
$\phi=\alpha-\arccos\left(\mathcal{C}\right)\hskip 11.38109pt\text{with}\hskip 11.38109pt\mathcal{C}=-\frac{\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}+\cos{\beta}\cos{\Theta}}{\sin{\beta}\sin{\Theta}}\,.$ (85)
The arccos can have two signs, so as in the two-dimensional case, two solutions exist. We found that when starting with the flat surface configuration, a minus sign has to be chosen. However, different from the two-dimensional case, the alternative solution (++ instead of -) also represents a valid Totimorphic configuration, although the two obtained configurations are far apart. In the case $\sin{\beta}=0$, $\phi$ can be freely chosen and $\Theta$ has a unique solution $\Theta=\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}\cos{\beta}}\right)$. In this case, the spring vector $\boldsymbol{s}^{-}$ is parallel to the z-axis. Any solution given by $\Theta$ will now form a triangle, from $\mathrm{A^{\prime}}$ to $\mathrm{P^{\prime}}$, from $\mathrm{P^{\prime}}$ to $\mathrm{C}$, and from $\mathrm{C}$ back to $\mathrm{A^{\prime}}$ (which is parallel to the z-axis). Changing $\phi$ will now rotate this triangle around the z-axis without distorting it, hence there is an infinite amount of valid configurations that lie on a circle around the z-axis.
$1\leq\mathcal{C}\leq 1\,.$ (86)
$-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\geq\cos{\beta}\cos{\Theta}-\sin{\beta}\sin{\Theta}\,.$ (87)
$-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\geq\cos\left(2\pi-\beta-\Theta\right)\,.$ (88)
which yields
$\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\right)-\beta\leq\Theta\leq 2\pi-\beta-\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\right)\,.$ (89)
$\beta-\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\right)\leq\Theta\leq\beta+\arccos\left(-\frac{\|\boldsymbol{s}^{-}\|}{l_{\text{b}}}\right)\,.$ (90)
4.3.4 Proof: deriving allowed range of $\theta_{z}$
Finally, we can calculate an allowed range for $\theta_{z}$ (as in the 2D case). This can be derived by requiring that the $\arccos$ in Eq. 89 and 90 has a valid solution, i.e.,
$\|\boldsymbol{s}^{-}\|^{2}\leq l_{\text{b}}^{2}\,.$ (91)
First, we rewrite $\boldsymbol{s}^{-}$ by expanding it:
$\boldsymbol{s}^{-}=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}=\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{P}}-\frac{l_{\text{b}}}{2}\cdot R_{y}(\varphi_{y})R_{z}(\varphi_{z})R_{x}(\theta_{x})R_{z}(\theta_{z})\mathbb{1}_{x}$ (92)
$R_{x}(-\theta_{x})R_{z}(-\varphi_{z})R_{y}(-\varphi_{y})\boldsymbol{s}^{-}=\boldsymbol{k}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\xi}^{\theta_{z},\frac{\pi}{2}}\,$, (93)
with $\boldsymbol{k}=R_{x}(-\theta_{x})R_{z}(-\varphi_{z})R_{y}(-\varphi_{y})\cdot\left(\boldsymbol{r}^{\mathrm{A^{\prime}}}-\boldsymbol{r}^{\mathrm{P}}\right)$. Since rotations leave the length of a vector unchanged, we can thus rewrite condition Eq. 91 as
$\|\boldsymbol{k}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{\xi}^{\theta_{z},\frac{\pi}{2}}\|^{2}\leq l_{\text{b}}^{2}\,.$ (94)
We solve this the same way as in Eq. 58, resulting in
$\boldsymbol{\xi}^{\theta_{k},\phi_{k}}\cdot\boldsymbol{\xi}^{\theta_{z},\frac{\pi}{2}}\geq\frac{\|\boldsymbol{k}\|}{l_{\text{b}}}-\frac{3}{4}\frac{l_{\text{b}}}{\|\boldsymbol{k}\|}\,$, (95)
$\cos\left(\theta_{k}-\theta_{z}\right)\geq C\hskip 11.38109pt\text{with}\hskip 11.38109ptC=\frac{\|\boldsymbol{k}\|}{l_{\text{b}}\sin{\phi_{k}}}-\frac{3}{4}\frac{l_{\text{b}}}{\|\boldsymbol{k}\|\sin{\phi_{k}}}\,$, (96)
from which the final interval is derived
$\theta_{k}-\arccos{C}\leq\theta_{z}\leq\theta_{k}+\arccos{C}\,.$ (97)
4.3.5 Building a three-dimensional Totimorphic structure